# Current address: ]College of William and Mary, Williamsburg, Virginia 23187-8795

### v1

@@ V.6 Extraction of $g_{1}$ and $A_{1}F_{1}$
  • The asymmetry $A_{1}$ in the DIS region was based on a fit to world data on inclusive spin structure functions in the DIS region (COMPASS, SLAC, HERMES, and JLab early 2010’s). The measured $A_{1}(x,Q^{2})$ values were fit, with simple Regge (powers of $x$) behavior assumed as $x\to 0$. The full uncertainty matrix of the fit was used to evaluate the systematic uncertainties.
  • The asymmetry $A_{2}$ in the DIS region was determined from $g_{2}^{WW}$ [118]. To evaluate systematic uncertainties, we added an extra twist-3 term obtained by fitting to the SLAC E155x data.
@@ V.9 Summary of Systematic Uncertainties
Figure 10: [Color online] Results on $g_{1}$ of the proton (red solid circles) plotted vs. $x$, compared with our standard parametrization (blue curves) and results from EG1b (black crossbars) [85]. The error bars are statistical only. The total systematic uncertainty is shown as the green band on the bottom edge of each panel and includes both experimental systematics and that from models used in the analysis. The two vertical lines in each panel indicate the location of the $\Delta(1232)$ and the integration limit $x_{\mathrm{hi}}$ that corresponds to $W=1.15$ GeV, see Section VII-A. Note that the $Q^{2}$ bins (in GeV2) are ordered column-wise, first from top to bottom, then from left to right.
@@
    Model uncertainties, analyzed by using the model variations described in Section V-G.
@@
    Systematic uncertainty due to the radiative tail from proton elastic scattering and from radiative corrections included in our Monte Carlo simulation. There are two potential contributions to this uncertainty: the models for the form factors and structure functions that enter these tails and radiative corrections, and the amount of target material traversed by the beam, which affects the external radiative tails. The former are “automatically” accounted for by our study of the overall systematic effect of using different structure function models, see Section V-G, and by trying different form factor parametrizations (see Section V-D). Note that we modified those model inputs simultaneously in the entire analysis chain, and retained the sign and magnitude of all variations to propagate them to the final results for the structure functions and moments, to properly account for correlations. The second uncertainty comes mostly from the uncertainty of the overall areal density of the target material, which in turn is determined by the packing fraction. The packing fraction (PF) is defined as the ratio of the actual quantity of ammonia target material contained in the target cell to the ideal quantity for a complete filling of the cell volume. While PF is not explicit in Eq. (39), it affects the total target thickness and thus the radiative effects in the simulation. To obtain the uncertainty due to PF, we re-ran the full simulation using a set of “alternate PF values” that differ from the nominal PF values. The data were re-analyzed using the alternate PF simulations and $g_{1}$ and $A_{1}F_{1}$ were extracted. The differences in the structure function results and those obtained using nominal PF values were taken as the uncertainty due to PF values. The uncertainty in the PF values themselves is taken to be 0.1. The nominal PF values used in the analysis are shown in Table III.
@@
    Systematic uncertainty due to normalizing the elastic peak of simulation to data. This includes: an uncertainty of $(1+E_{b})\%$ (with $E_{b}$ the beam energy in GeV) (see Section V-D) and the statistical uncertainty of the data elastic or quasi-elastic peak (Fig. 9).
@@
    Systematic uncertainty due to CC efficiency. This was obtained by re-running the analysis without applying the CC efficiency to the simulated spectra. The difference in the extracted $g_{1}$ and $A_{1}F_{1}$ from the nominal analysis indicates the effect of the CC efficiency, and 20% of this difference was taken as the systematic uncertainty due to the CC efficiency. Details of CC efficiency analysis were described in Section IV-C.
@@
    Systematic uncertainty due to background contributions. This includes effect from pion and pair production background (1% of the polarized yield, see Section IV-D); 15N elastic tail in the inelastic region (see Section V-H); and the 15N inelastic scattering background in the inelastic polarized yield, which contributes 0.7% of the polarized yield estimated by calculating the 15N to proton inelastic polarized cross section ratio and applying the polarization ratio of the two.
@@
V.10 Results on neutron $g_{1}$
Figure 11: [Color online] Results on $g_{1}$ of the neutron (red solid circles) vs. $x$ extracted from deuteron and proton data. These are compared with our standard parametrization with the best $A_{1}$ model (blue curves) and results from EG1b (black crossbars) [85]. The error bars are statistical only. The total systematic uncertainty is shown as the green band on the bottom edge of each panel that includes both experimental systematics and that from models used in the analysis.
From our results on $g_{1}$ of the proton and the deuteron, we further extracted values for the neutron spin structure function, $g_{1}^{n}$, which were not available previously. This extraction is based on the simplified model of the deuteron structure functions described in Section II-E, where the structure functions of the deuteron are expressed as sum of the two nucleon structure functions, convoluted with their spin and momentum distributions in the deuteron [90]. To extract the deconvoluted structure functions for a neutron at rest, we followed a similar procedure as for the extraction of $g_{1}^{p,d}$ from the data on the proton and deuteron, described in detail below.
$g_{1}^{n,\mathrm{(data,free)}}$ $=$ $g_{1}^{n,\mathrm{model,free}}$
$+(g_{1}^{n,\mathrm{model-a(b),free}}-g_{1}^{n,\mathrm{model,free}})$
Note that the two versions ($a,b$) of the modification produce nearly identical results. The statistical and experimental systematic uncertainties of each of the input $g_{1}^{p}$ and $g_{1}^{d}$ were propagated to produce the experimental uncertainties in the extracted $g_{1}^{n}$. The uncertainty in this extraction was studied using a similar method as in Section V-G, but now by changing the initial model for the neutron.
Figure 12: [Color online] Results on the isovector component $g_{1}^{p-n}$. See Fig. 11 caption for details.
Finally, we show the isovector combination $g_{1}^{p-n}(x,Q^{2})$ in Fig. 12. Comparing with Figs. 10, 11, and 13, one clearly observes the suppression in the isovector component of the $\Delta^{(1232)}$. This has important consequences for the Bjorken integral (i.e., the first moment of $g_{1}^{p-n}(x,Q^{2})$), see Section VII-F, and its comparison to $\chi$EFT [121]. Like the other EG4 results on $g_{1}$ and $A_{1}F_{1}$, Fig. 12 also makes clear that in the $(x,Q^{2})$ domain covered by EG4, the parameterization of the world data (continuous blue line) needs to be revised.
@@ VI.1 Results on $g_{1}$ and $A_{1}F_{1}$ for the Proton and Deuteron
Figure 13: [Color online] Results on $g_{1}$ of the deuteron. See Fig. 10 caption for details.
@@
Figure 14: [Color online] Results on $A_{1}F_{1}$ of the proton plotted vs. $W$ to emphasize the contribution of different resonance regions. Note that earlier CLAS data [85] are not available for this quantity. See Fig. 10 caption for more details.
Figure 15: [Color online] Results on $A_{1}F_{1}$ of the deuteron plotted vs. $W$ to emphasize the contribution of different resonance regions. Note that earlier CLAS data [85] are not available for this quantity. See Fig. 10 caption for more details.
Our final results for $g_{1}$ and $A_{1}F_{1}$ of the proton and the deuteron are shown in Figs. 10 through 15. Our results are consistent with the previous data [86, 85] where their coverage overlap, but extend the measured $Q^{2}$ range to three times smaller values, below the pion mass squared ($m_{\pi}^{2}$). This makes it possible to rigorously test $\chi$EFT calculations for the nucleon spin structure functions. Comparing to our existing $A_{1}$ model, which was based on existing data prior to EG4, one can see that it tends to be more positive than the data points, especially at lower $Q^{2}$. Our new data can help improve this model.
@@ VII.1 General Procedure Used to Form the Integrals
where $x_{\mathrm{lo}}$ and $x_{\mathrm{hi}}$ are the lower and upper limits of the data integration, $g_{1}^{\mathrm{mod}}$ is the estimate from our best model described in Section V-F and $g_{1}^{\mathrm{EG4~{}data}}$ is our data. The low-$x$ integration starts at $x=10^{-3}$ and the step size was set to $dx=10^{-5}$.
The range $(x_{\mathrm{lo}},x_{\mathrm{hi}})$ of the data integration in Eq. (56) is slightly narrower than the actual kinematic coverage of the experiment because the data uncertainties on the edge of the coverage are large: At large $x$, near the pion threshold $W\approx 1.07$ GeV, the polarized yield is dominated by the radiative tail from elastic scattering and the measured polarized yield is not sensitive to the value of $g_{1}$ or $A_{1}F_{1}$; while for the low $x$ or large $W$ edge of the data, the spectrometer acceptance drops and the statistics of the data is low, see Section V-I. Thus, the value of $x_{\mathrm{hi}}$ was chosen to be at $W=1.15$ GeV and $x_{\mathrm{lo}}$ was set at a few tens of MeV below the maximum $W$ covered by the data. The upper $x$-limits of the data integration ranges are shown in Figs. 10 and 13.
Uncertainties in the model enters the integral in two ways. First, the model enters the extraction of the data itself, as described in Section V-G. Second, the uncertainty of the model integration below $x_{\mathrm{lo}}$ and above $x_{\mathrm{hi}}$ was evaluated by variation of the model parameters, similar to the method described in Section V-G. When evaluating model uncertainties of the data integral, the sign of each model uncertainty of $g_{1}$ or $A_{1}F_{1}$ is retained to allow partial cancellation when integrating. The total uncertainty on the full integral was formed by adding all experimental and model uncertainties in quadrature.
@@ VII.2 Latest $\chi$EFT Moment Predictions
In the following sections, we will compare the EG4 result on the moments to theoretical predictions, in particular from $\chi$EFT. We will display only two recent predictions, from Ref. [53] and Ref. [54], as they supersede earlier $\chi$EFT calculations [122, 73, 75, 77]. The crucial advance of the latest calculations is that they account rigorously for the $\Delta^{(1232)}$ excitation. At the same time, the approaches by Refs. [53] and [54] have significant differences. The chief reason is the effective accounting of higher-order contributions of the $\chi$EFT series by Ref. [54], which uses a phenomenological form factor. It softens the steep $Q^{2}$-dependence otherwise seen in Ref. [53]. The softening form factor of Ref. [54] accounts for their typically better description of the $Q^{2}$-dependence of moments. The methodological differences between the two calculations are detailed in Section VII-G.
@@ VII.3 Moments of the Proton Spin Structure functions
Figure 16: [Color online] EG4 results on the proton $\overline{\Gamma}_{1}^{p}(Q^{2})$. Integrals over the full (experimentally covered) $x$-range are shown as solid red (open blue) circles. The inner and the outer error bars (sometimes too small to be seen) are for statistical and total uncertainties, respectively. To illustrate their relative contributions, we also show the systematic uncertainty as the (red colored) band at the bottom of the figure, which includes both experimental systematic and model uncertainties. These results are compared with earlier data from CLAS [85] (solid black crossbars), $\chi$EFT predictions by Alarcón et al. [54] (green backslash hatched band) and Bernard et al. [53] (magenta forward-slash hatched band), phenomenological models by Burkert and Ioffe [47, 48] (orange dot-dashed curve) and Pasechnik et al. [50] (grey dashed curve), and our own spin structure function parameterization [85] (purple solid curve). The MAID model [88] that includes only one-pion production contributions is shown by the (yellow) dotted curve. The GDH slope (black dotted line) is the GDH sum rule expectation for $Q^{2}\to 0$. See texts for detailed discussions.
@@
Figure 17: [Color online] EG4 results on the proton $\overline{I}_{TT}^{p}(Q^{2})$. The GDH value is shown as the short horizontal line, at $\overline{I}_{TT}^{p}(0)=-0.804$. The result from the photoproduction experiment [59, 61, 62, 60] is shown as the solid purple star at $Q^{2}=0$. Note that earlier CLAS data [85] and the phenomenological model predictions [48, 50] are not available for this integral. See Fig. 16 caption for more details.
@@
Figure 18: [Color online] EG4 results on the proton ${\overline{\gamma}_{0}}^{p}(Q^{2})$. The photoproduction data point [123] is slightly displaced from $Q^{2}=0$ (purple star). Note that phenomenological model predictions [48, 50] are not available for this integral. See Fig. 16 caption for more details.

### v2

@@ V.6 Extraction of $g_{1}$ and $A_{1}F_{1}$
  • The asymmetry $A_{1}$ in the DIS region was based on a fit to world data on inclusive spin structure functions in the DIS region (COMPASS, SLAC, HERMES, and JLab early 2010’s) [32, 33, 34, 35, 36, 37, 145, 146, 105, 106, 45, 46, 48, 47, 108, 50, 38, 89, 102, 98, 103, 147]. The measured $A_{1}(x,Q^{2})$ values were fit, with simple Regge (powers of $x$) behavior assumed as $x\to 0$. The full uncertainty matrix of the fit was used to evaluate the systematic uncertainties.
  • The asymmetry $A_{2}$ in the DIS region was determined from $g_{2}^{WW}$ [148]. To evaluate systematic uncertainties, we added an extra twist-3 term obtained by fitting to the SLAC E155x [149] data.
@@ V.9 Summary of Systematic Uncertainties
Figure 10: [Color online] Results on $g_{1}$ of the proton (red solid circles) plotted vs. $x$, compared with our standard parametrization (blue curves) and results from EG1b (black crossbars) [102]. The error bars are statistical only. The total systematic uncertainty is shown as the green band on the bottom edge of each panel and includes both experimental systematics and that from models used in the analysis. The size of the total systematic uncertainty varies from $0.013-0.15$ to $0.0045-0.19$, and the relative size of the total systematic to statistical uncertainties varies from $(8.0-33)\%$ to $(16-140)\%$, from the lowest to the highest $Q^{2}$ bins. The two vertical lines in each panel indicate the location of the $\Delta(1232)$ and the integration limit $x_{\mathrm{hi}}$ that corresponds to $W=1.15$ GeV, see Section VII.1. Note that the $Q^{2}$ bins (in GeV2) are ordered column-wise, first from top to bottom, then from left to right.
@@
    Model uncertainties, analyzed by using the model variations described in Section V.7.
@@
    Systematic uncertainty due to the radiative tail from proton elastic scattering and from radiative corrections included in our Monte Carlo simulation. There are two potential contributions to this uncertainty: the models for the form factors and structure functions that enter these tails and radiative corrections, and the amount of target material traversed by the beam, which affects the external radiative tails. The former are “automatically” accounted for by our study of the overall systematic effect of using different structure function models, see Section V.7, and by trying different form factor parametrizations (see Section V.4). Note that we modified those model inputs simultaneously in the entire analysis chain, and retained the sign and magnitude of all variations to propagate them to the final results for the structure functions and moments, to properly account for correlations. The second uncertainty comes mostly from the uncertainty of the overall areal density of the target material, which in turn is determined by the packing fraction. The packing fraction (PF) is defined as the ratio of the actual quantity of ammonia target material contained in the target cell to the ideal quantity for a complete filling of the cell volume. While PF is not explicit in Eq. (39), it affects the total target thickness and thus the radiative effects in the simulation. To obtain the uncertainty due to PF, we re-ran the full simulation using a set of “alternate PF values” that differ from the nominal PF values. The data were re-analyzed using the alternate PF simulations and $g_{1}$ and $A_{1}F_{1}$ were extracted. The differences in the structure function results and those obtained using nominal PF values were taken as the uncertainty due to PF values. The uncertainty in the PF values themselves is taken to be 0.1. The nominal PF values used in the analysis are shown in Table 3.
@@
    Systematic uncertainty due to normalizing the elastic peak of simulation to data. This includes: an uncertainty of $(1+E_{b})\%$ (with $E_{b}$ the beam energy in GeV) (see Section V.4) and the statistical uncertainty of the data elastic or quasi-elastic peak (Fig. 9).
@@
    Systematic uncertainty due to CC efficiency. This was obtained by re-running the analysis without applying the CC efficiency to the simulated spectra. The difference in the extracted $g_{1}$ and $A_{1}F_{1}$ from the nominal analysis indicates the effect of the CC efficiency, and 20% of this difference was taken as the systematic uncertainty due to the CC efficiency. Details of CC efficiency analysis were described in Section IV.3.
@@
    Systematic uncertainty due to background contributions. This includes effect from pion and pair production background (1% of the polarized yield, see Section IV.4); 15N elastic tail in the inelastic region (see Section V.8); and the 15N inelastic scattering background in the inelastic polarized yield, which contributes 0.7% of the polarized yield estimated by calculating the 15N to proton inelastic polarized cross section ratio and applying the polarization ratio of the two.
@@ VI.1 Results on $g_{1}$ and $A_{1}F_{1}$ for the Proton and Deuteron
Figure 11: [Color online] Results on $g_{1}$ of the deuteron. The size of the total systematic uncertainty varies from $0.020-0.183$ to $0.010-0.294$, and the relative size of the systematic to statistical uncertainty varies from $(4.3-52)\%$ to $(16-55)\%$, from the lowest to the highest $Q^{2}$ bins. See Fig. 10 caption for more details.
@@
Figure 12: [Color online] Results on $A_{1}F_{1}$ of the proton plotted vs. $W$ to emphasize the contribution of different resonance regions. The size of the total systematic uncertainty varies from $0.0134-0.151$ to $0.0045-0.190$, and the relative size of the systematic to statistical uncertainty varies from $(8.2-33)\%$ to $(8.0-118)\%$, from the lowest to the highest $Q^{2}$ bins. Note that earlier CLAS data [102] are not available for this quantity. See Fig. 10 caption for more details.
Figure 13: [Color online] Results on $A_{1}F_{1}$ of the deuteron plotted vs. $W$ to emphasize the contribution of different resonance regions. The size of the total systematic uncertainty varies from $0.020-0.22$ to $0.019-0.952$, and the relative size of the systematic to statistical uncertainty varies from $(4.3-55)\%$ in the lowest Q2 bin, to $(18-39)\%$, from the lowest to the highest $Q^{2}$ bins. Note that earlier CLAS data [102] are not available for this quantity. See Fig. 10 caption for more details.
Our final results for $g_{1}$ and $A_{1}F_{1}$ of the proton and the deuteron are shown in Figs. 10 through 13. Our results are consistent with the previous data [103, 102] where their coverage overlap, but extend the measured $Q^{2}$ range to three times smaller values, below the pion mass squared ($m_{\pi}^{2}$). This makes it possible to rigorously test $\chi$EFT calculations for the nucleon spin structure functions. Comparing to our existing $A_{1}$ model, which was based on existing data prior to EG4, one can see that it tends to be more positive than the data points, especially at lower $Q^{2}$. Our new data can help improve this model.
VI.2 Results on neutron $g_{1}$
Figure 14: [Color online] Results on $g_{1}$ of the neutron (red solid circles) vs. $x$ extracted from deuteron and proton data. These are compared with our standard parametrization with the best $A_{1}$ model (blue curves) and results from EG1b (black crossbars) [102]. The error bars are statistical only. The total systematic uncertainty is shown as the green band on the bottom edge of each panel that includes both experimental systematics and that from models used in the analysis. The size of the total systematic uncertainty varies from $0.040-0.305$ to $0.023-0.160$, and the relative size of the systematic to statistical uncertainty varies from $(10.7-55)\%$ to $(32-170)\%$, from the lowest to the highest $Q^{2}$ bins.
From our results on $g_{1}$ of the proton and the deuteron, we further extracted values for the neutron spin structure function, $g_{1}^{n}$, which were not available previously. This extraction is based on the simplified model of the deuteron structure functions described in Section II.5, where the structure functions of the deuteron are expressed as sum of the two nucleon structure functions, convoluted with their spin and momentum distributions in the deuteron [115]. To extract the deconvoluted structure functions for a neutron at rest, we followed a similar procedure as for the extraction of $g_{1}^{p,d}$ from the data on the proton and deuteron, described in detail below.
$g_{1}^{n,\mathrm{(data,free)}}$ $=$ $g_{1}^{n,\mathrm{model,free}}$
$+(g_{1}^{n,\mathrm{model-a(b),free}}-g_{1}^{n,\mathrm{model,free}})$
Note that the two versions ($a,b$) of the modification produce nearly identical results. The statistical and experimental systematic uncertainties of each of the input $g_{1}^{p}$ and $g_{1}^{d}$ were propagated to produce the experimental uncertainties in the extracted $g_{1}^{n}$. The uncertainty in this extraction was studied using a similar method as in Section V.7, but now by changing the initial model for the neutron.
Figure 15: [Color online] Results on the isovector component $g_{1}^{p-n}$. Across all $Q^{2}$ bins, the size of the total systematic uncertainty varies from $0.022$ to $1.013$, while the relative size of the total systematic to statistical uncertainty varies within $(7.7-173)\%$. See Fig. 14 caption for more details.
Finally, we show the isovector combination $g_{1}^{p-n}(x,Q^{2})$ in Fig. 15. Comparing with Figs.10, 11, and 14, one clearly observes the suppression in the isovector component of the $\Delta^{(1232)}$. This has important consequences for the Bjorken integral (i.e., the first moment of $g_{1}^{p-n}(x,Q^{2})$), see Section VII.6, and its comparison to $\chi$EFT [151]. Like the other EG4 results on $g_{1}$ and $A_{1}F_{1}$, Fig. 15 also makes clear that in the $(x,Q^{2})$ domain covered by EG4, the parameterization of the world data (continuous blue line) needs to be revised.
@@ VII.1 General Procedure Used to Form the Integrals
where $x_{\mathrm{lo}}$ and $x_{\mathrm{hi}}$ are the lower and upper limits of the data integration, $g_{1}^{\mathrm{mod}}$ is the estimate from our best model described in Section V.6 and $g_{1}^{\mathrm{EG4~{}data}}$ is our data. The low-$x$ integration starts at $x=10^{-3}$ and the step size was set to $dx=10^{-5}$.
The range $(x_{\mathrm{lo}},x_{\mathrm{hi}})$ of the data integration in Eq. (56) is slightly narrower than the actual kinematic coverage of the experiment because the data uncertainties on the edge of the coverage are large: At large $x$, near the pion threshold $W\approx 1.07$ GeV, the polarized yield is dominated by the radiative tail from elastic scattering and the measured polarized yield is not sensitive to the value of $g_{1}$ or $A_{1}F_{1}$; while for the low $x$ or large $W$ edge of the data, the spectrometer acceptance drops and the statistics of the data is low, see Section V.9. Thus, the value of $x_{\mathrm{hi}}$ was chosen to be at $W=1.15$ GeV and $x_{\mathrm{lo}}$ was set at a few tens of MeV below the maximum $W$ covered by the data. The upper $x$-limits of the data integration ranges are shown in Figs. 10 and 11.
Uncertainties in the model enters the integral in two ways. First, the model enters the extraction of the data itself, as described in Section V.7. Second, the uncertainty of the model integration below $x_{\mathrm{lo}}$ and above $x_{\mathrm{hi}}$ was evaluated by variation of the model parameters, similar to the method described in Section V.7. When evaluating model uncertainties of the data integral, the sign of each model uncertainty of $g_{1}$ or $A_{1}F_{1}$ is retained to allow partial cancellation when integrating. The total uncertainty on the full integral was formed by adding all experimental and model uncertainties in quadrature.
@@ VII.2 Latest $\chi$EFT Moment Predictions
In the following sections, we will compare the EG4 result on the moments to theoretical predictions, in particular from $\chi$EFT. We will display only two recent predictions, from Ref. [59] and Ref. [60], as they supersede earlier $\chi$EFT calculations [152, 79, 81, 83]. The crucial advance of the latest calculations is that they account rigorously for the $\Delta^{(1232)}$ excitation. At the same time, the approaches by Refs. [59] and [60] have significant differences. The chief reason is the effective accounting of higher-order contributions of the $\chi$EFT series by Ref. [60], which uses a phenomenological form factor. It softens the steep $Q^{2}$-dependence otherwise seen in Ref. [59]. The softening form factor of Ref. [60] accounts for their typically better description of the $Q^{2}$-dependence of moments. The methodological differences between the two calculations are detailed in Section VII.7.
@@ VII.3 Moments of the Proton Spin Structure functions
Figure 16: [Color online] EG4 results on the proton $\overline{\Gamma}_{1}^{p}(Q^{2})$. Integrals over the full (experimentally covered) $x$-range are shown as solid red (open blue) circles. The inner and the outer error bars (sometimes too small to be seen) are for statistical and total uncertainties, respectively. To illustrate their relative contributions, we also show the systematic uncertainty as the (red colored) band at the bottom of the figure, which includes both experimental systematic and model uncertainties. Across all $Q^{2}$ bins, the size of the systematic uncertainty varies from $0.0002$ to $0.0068$, while the relative size of the systematic to statistical uncertainty varies within $(23-233)\%$. These results are compared with earlier data from CLAS [102] (solid black crossbars), $\chi$EFT predictions by Alarcón et al. [60] (green backslash hatched band) and Bernard et al. [59] (magenta forward-slash hatched band), phenomenological models by Burkert and Ioffe [53, 54] (orange dot-dashed curve) and Pasechnik et al. [56] (grey dashed curve), and our own spin structure function parameterization [102] (purple solid curve). The MAID model [113] that includes only one-pion production contributions is shown by the (yellow) dotted curve. The GDH slope (black dotted line) is the GDH sum rule expectation for $Q^{2}\to 0$. See texts for detailed discussions.
Figure 17: [Color online] EG4 results on the proton $\overline{I}_{TT}^{p}(Q^{2})$. Across all $Q^{2}$ bins, the size of the systematic uncertainty (red colored band) varies from $0.186$ to $0.104$, while the relative size of the systematic to statistical uncertainty varies within $(54-357)\%$. The GDH value is shown as the short horizontal line, at $\overline{I}_{TT}^{p}(0)=-0.804$. The result from the photoproduction experiment [65, 67, 68, 66] is shown as the solid purple star at $Q^{2}=0$. Note that earlier CLAS data [102] and the phenomenological model predictions [54, 56] are not available for this integral. See Fig. 16 caption for more details.
@@
Figure 18: [Color online] EG4 results on the proton ${\overline{\gamma}_{0}}^{p}(Q^{2})$. Across all $Q^{2}$ bins, the size of the systematic uncertainty (red colored band) varies from $0.0045$ to $0.198$, while the relative size of the systematic to statistical uncertainty varies within $(66-364)\%$. The photoproduction data point [153] is slightly displaced from $Q^{2}=0$ (purple star). Note that phenomenological model predictions [54, 56] are not available for this integral. See Fig. 16 caption for more details.